# Hydrodynamics of Relativistic Superheated Bubbles

### v1

@@ Abstract
Relativistic, charged, superheated bubbles may play an important role in neutron star mergers if first-order phase transitions are present in the phase diagram of Quantum Chromodynamics. We describe the properties of these bubbles in the hydrodynamic regime. We find two qualitative differences with supercooled bubbles. First, the pressure inside an expanding superheated bubble can be higher or lower than the pressure outside the bubble. Second, some fluid flows develop metastable regions behind the bubble wall. We consider the possible role of a conserved charge akin to baryon number. The fluid flow profiles are unaffected by this charge if the speed of sound is constant in each phase, but they are modified for more general equations of state. We compute the efficiency factor relevant for gravitational wave production.
CPHT-RR050.062024
1 Introduction
@@
Figure 1: Two possible phase transitions in QCD, indicated by the solid red curves. $T$ and $\mu$ are the temperature and the baryon chemical potential, respectively. The dotted black curve on the left panel shows a possible evolution of a region of a NS merger as this region is heated and/or compressed. The points dubbed $A$, $A^{\prime}$ and $C$ correspond to the states shown in Fig. 2. See text and Ref. Casalderrey-Solana:2022rrn .
@@
The energy density along the black dotted curve in Fig. 1 displays the multivalued form characteristic of a FOPT, as shown in Fig. 2. The superheated region is pushed from left to right along the lower branch of the phase diagram. Once the superheating is large enough, namely once the region of interest is sufficiently deep into the metastable branch, bubbles of the stable phase begin to nucleate. The point where this happens is labeled “$B$” in Fig. 2, the nucleated phase is labelled “$C$”, and the transition is indicated by a vertical, solid, black arrow. We will refer to these nucleated bubbles as “superheated bubbles”. The dynamics of these bubbles can produce gravitational waves that would provide direct experimental access to the QCD phase transition Casalderrey-Solana:2022rrn . The order-of-magnitude estimate in this reference shows that the frequency of these gravitational waves falls roughly in the MHz range, and that they may be potentially observable with future detectors.
Figure 2: Energy density as a function of the position along the black dotted curve in Fig. 1. Both $T$ and $\mu$ increase from left to right. The black, dotted, vertical line indicates the location of the phase transition, determined by the condition that the states $A$ and $A^{\prime}$ have the same free energy density. The green and blue curves indicate stable and metastable states, respectively. As some region of the NS merger is sufficiently heated and/or compressed, it enters the lower metastable branch. At the point $B$ bubbles of the preferred state $C$ on the upper stable branch are nucleated. The direction of this phase transition is the opposite of that in a supercooled phase transition, which would take place as indicated by the vertical, dotted, red arrow.
@@
The dynamics of relativistic supercooled bubbles has been extensively studied in the literature, motivated by their potential implications for cosmological FOPT — see e.g. Hindmarsh:2020hop for a review. Motivated by their possible role in NS mergers, in this paper we initiate the study of relativistic, charged, superheated bubbles, focusing on their hydrodynamic properties in the simple case of a bag-model, conformal EoS. We find two qualitative differences compared to supercooled bubbles. First, the pressure inside a superheated bubble may be higher or lower than the pressure outside the bubble. In contrast, in supercooled bubbles the inside pressure is always higher than the outside pressure. Second, some fluid flows develop regions behind the bubble wall that are metastable for any choice of parameters in the EoS, which would therefore decay at sufficiently long times. In contrast, in supercooled transitions there always exists a choice of parameters for which the flow behind the wall only explores stable states.
@@
Part of the contents of this paper overlap with those in Ref. Barni:2024lkj , which appeared while this paper was being typewritten.
2 Self-similar profiles
@@
[…]
@@
If each branch of the EoS is conformal, namely, if the pressure in each phase takes the form
$p(T,\mu)=T^{d+1}f\left(\frac{\mu}{T}\right)+\mbox{const.}\,,$. , (11)
with $f$ an arbitrary function, then the speed of sound is constant $c_{s}^{2}=1/d$. In this case (8) becomes a nested system that can be sequentially solved for $v$, $w$ and $n$, similarly to what happens in the neutral case Espinosa:2010hh . The nested structure means that the addition of charge does not change the properties of self-similar flows for conformal EoS. In fact, the last equation simply implies that $n\propto w^{1/(1+c_{s}^{2})}$.
In contrast, if the EoS is non-conformal, the speed of sound will depend on the local properties of the plasma, $c_{s}=c_{s}(e,n)$, making equations (8) coupled to one another. In this case the effect of the charge is already present at the level of the self-similar flow. We leave studies of non-conformal theories for future work.
@@
where $+$(--) refers to the plasma in front (behind) the discontinuity, the velocities $v_{\pm}$ are measured in the rest frame of the discontinuity, and $\gamma$ is the usual Lorentz factor. Rearranging the first two matching conditions we arrive at
@@
Eqs. (8) and (13) determine the entire fluid flow as a function of three input parameters: the wall velocity $\xi_{w}$, the nucleation energy density $e_{n}$ and the nucleation charge density $n_{n}$. In the following we will solve these equations in the case of spherical bubbles in three spatial dimensions ($d=3$) and a constant speed of sound.
When $c_{s}$ is constant we can solve for the velocity in (8) independently of the rest of the variables. By assuming different initial conditions $v(\xi_{0})=v_{0}$ one can obtain all possible flows that are self-similar, shown in Fig. 3. Contrary to the supercooled case, we seek solutions where at $\xi>0$ (expanding bubbles) the flow has negative speed $v<0$ (energy inflow). In order to produce Fig. 3, we chose $c_{s}^{2}=1/3$. We expect that different values of the speed of sound will change the picture quantitatively but not qualitatively.
@@
There is yet another kind of flows, those that are contained in the shaded area of Fig. 3. Given that they would become multivalued, these flows cannot cross the dot-dashed line, and they must follow to higher $\xi$. As they cannot smoothly connect to a state with $v=0$, they jump discontinously to it at the shock location. This shock is different to the one present in hybrids and detonations as the flow in motion is behind and not ahead. An entropic analysis shows that the shock front in Fig. 3 increases entropy only if the flow in motion is ahead. Hence, no flow starting in the mentioned area is realizable.
@@
3 Solutions for a conformal equation of state
In order to analyze the fluid flows in more detail, a choice of an EoS is needed. Assuming two branches such that $c_{s}^{2}=1/3$ in each of them, we choose
$p_{H}(T,\mu)$ $=T^{4}\left(a_{H}+b_{H}\frac{\mu^{2}}{T^{2}}+c_{H}\frac{\mu^{4}}{T^{4}}\right)-\epsilon\,,$ (14)
$p_{L}(T,\mu)$ $=T^{4}\left(a_{L}+b_{L}\frac{\mu^{2}}{T^{2}}+c_{L}\frac{\mu^{4}}{T^{4}}\right)\,,$
@@ 3 Solutions for a conformal equation of state
where $a_{H,L}$, $b_{H,L}$ and $c_{H,L}$ reflect the number of degrees of freedom in each phase and “$H$” and “$L$” refer to the high- and the low-energy phases, respectively. The constant $\epsilon$ is sometimes referred to as the “bag constant” and it measures the energy difference between the two phases in vacuum, that is, at $T=\mu=0$. For the choices $a_{H}=b_{H}=c_{H}=1$ and $a_{L}=b_{L}=c_{L}=1/2$, the phase diagram is shown in Fig. 4: There is a curve of FOPT transition at $T^{2}+\mu^{2}=\sqrt{2\epsilon}$. We do not choose this EoS because we expect it to provide a realistic description of QCD but because it is the simplest example in which an investigation of superheated bubbles is possible. We will report on more realistic examples elsewhere.
@@
Figure 4: Phase diagram for the choice $a_{H}=b_{H}=c_{H}=1$ and $a_{L}=b_{L}=c_{L}=1/2$. “$L$” and “$H$” refer to the low- and high-energy phases, respectively. All quantities are measured in units of the bag constant, $\epsilon$, which is the energy difference between the two phases at $T=\mu=0$.
@@
The specific detonation shown in Fig. 6(middle) has been chosen to illustrate that, for some flows, the pressure in a region just behind the wall, in this case in the region $\xi\lesssim 0.8$, can be negative. This is necessarily lower than the pressure of the $L$-phase, which is always positive regardless of the temperature and of the values of the parameters $a_{H,L}$, $b_{H,L}$ and $c_{H,L}$ that specify the EoS. This is in stark contrast with supercooled flows, for which there is always a choice of these parameters for which the flow behind the wall only explores stable states. Instead, in the superheated case, at arbitrarily late times there is an arbitrarily large, approximately homogeneous region behind the wall that is thermodynamically disfavoured with respect to the $L$-phase at the same temperature. This region will eventually decay by nucleating bubbles of the $L$-phase if the bubble is allowed to expand for arbitrarily long times. Whether this happens in practice will depend on the nucleation rates in front and behind the bubble wall, on the bubble wall velocities, etc. Nevertheless, from the viewpoint of the dynamics of a single bubble, flows with negative pressure may regarded as thermodynamically unstable. In Fig. 7 these flows correspond to points above the blue curve labelled as $p_{-}=0$. Note that deflagrations cannot suffer from this problem since the phase right behind the wall is directly the stable $H$-phase.
@@ 4 Bounds
Let us start with detonations. Given a wall velocity $\xi_{w}$, consider increasing the strength $\alpha_{n}$ or, equivalently, decreasing the nucleation energy $e_{n}$ relative to the bag constant $\epsilon$. This decreases the energy behind the wall $e_{-}$ and increases the speed $v_{-}$. Clearly, the strongest transition we can consider is that for which the energy density in the fluid behind the wall reaches its vacuum energy, that is, $e_{-}=\epsilon$. Precisely at this point we also reach the maximum possible velocity behind the wall, $v_{-}=1$. Solving the matching conditions with these values results in the bound
@@
We see that, as we increase the strength factor, the first solutions that cease to exist are the hybrid flows. The next ones to be forbidden are detonations. Therefore, for strong transitions only deflagrations survive, but they are confined to regions with very low wall velocities. This picture is opposite to that in the supercooled case (Espinosa:2010hh, ), where deflagrations cease to exist for strong transitions while hybrids and detonations become confined to higher and higher wall velocities.
@@
5 Efficiency factor
Ref. Casalderrey-Solana:2022rrn showed that the dynamics of superheated bubbles in NS mergers can give rise to an interesting GW signal in the MHz range. A relevant quantity to estimate the amount of GW production is the so called efficiency factor Hindmarsh:2020hop
[…]
which is a measure of how much energy has been deposited in the motion of the fluid. We have computed the efficiency factor for several strength factors as a function of the wall velocity in the allowed regions where we can find bubble solutions. The result is summarized in Fig. 8. The empty regions in this figure arise form the disallowed regions, above the solid black curves, in Fig. 7.
Figure 8: Efficiency factor (22) as a function of the bubble wall velocity for several choices of the strength factor.
The behavior of the efficiency factor for deflagrations and detonations is quite simple, since it is monotonous in both cases. For deflagrations the efficiency factor increases with the wall velocity at fixed strength, whereas for detonations it decreases. In both cases the efficiency factor increases with the strength factor at fixed wall velocity. Hybrids exhibit a richer, non-monotonic behaviour. For low $\alpha_{n}$, hybrids connect the deflagrations with the detonations and the efficiency factor has a maximum at some intermediate value of the wall velocity. For moderate strengths, the maximum efficiency factor increases with the strength factor. As the strength is increased, the low-velocity hybrids cease to exist and the efficiency factor becomes a monotonically decreasing function of the wall velocity. At this stage, increasing the strength implies discarding hybrid solutions and the maximum efficiency factor decreases. In summary, the maximum efficiency factor corresponds to a hybrid at some intermediate values of both the wall velocity and the strength factor.

### v2

# CPHT-RR050.062024

@@ Abstract
Relativistic, charged, superheated bubbles may play an important role in neutron star mergers if first-order phase transitions are present in the phase diagram of Quantum Chromodynamics. We describe the properties of these bubbles in the hydrodynamic regime. We find two qualitative differences with supercooled bubbles. First, the pressure at the center of an expanding superheated bubble can be higher or lower than the pressure in the asymptotic, metastable phase. Second, some fluid flows develop metastable regions behind the bubble wall for any choice of the equation of state. We consider the possible role of a conserved charge akin to baryon number. The fluid flow profiles are unaffected by this charge if the speed of sound is constant in each phase, but they are modified for more general equations of state. We compute the efficiency factor relevant for gravitational wave production.
I Introduction
@@
Figure 1: Two possible phase transitions in QCD, indicated by the solid red curves. $T$ and $\mu$ are the temperature and the baryon chemical potential, respectively. The dotted black curve on the left panel shows a possible evolution of a region of a NS merger as this region is heated and/or compressed. The points dubbed AA, $A^{\prime}$ and $C$ correspond to the states shown in Fig. 2. See text and Ref. Casalderrey-Solana:2022rrn .
@@
The energy density along the black dotted curve in Fig. 1 displays the multivalued form characteristic of a FOPT, as shown in Fig. 2. The superheated region is pushed from left to right along the lower branch of the phase diagram. Once the superheating is large enough, namely once the region of interest is sufficiently deep into the metastable branch, bubbles of the stable phase begin to nucleate. [^1] To avoid confusion, we clarify that by “bubble” we mean the volume of a given phase separated from the outside phase by the “bubble wall”, which is a surface. The point where this happens is labeled “BB” in Fig. 2, the nucleated phase is labelled “CC”, and the transition is indicated by a vertical, solid, black arrow. We will refer to these nucleated bubbles as “superheated bubbles”. The dynamics of these bubbles can produce gravitational waves that would provide direct experimental access to the QCD phase transition Casalderrey-Solana:2022rrn . The order-of-magnitude estimate in this reference shows that the frequency of these gravitational waves falls roughly in the MHz range, and that they may be potentially observable with future detectors.
Figure 2: Energy density as a function of the position along the black dotted curve in Fig. 1. Both $T$ and $\mu$ increase from left to right. The black, dotted, vertical line indicates the location of the phase transition, determined by the condition that the states AA and $A^{\prime}$ have the same free energy density. The green and blue curves indicate stable and metastable states, respectively. As some region of the NS merger is sufficiently heated and/or compressed, it enters the lower metastable branch. At the point $B$ bubbles of the preferred state $C$ on the upper stable branch are nucleated. The direction of this phase transition is the opposite of that in a supercooled phase transition, which would take place as indicated by the vertical, dotted, red arrow.
@@
The dynamics of relativistic supercooled bubbles has been extensively studied in the literature, motivated by their potential implications for cosmological FOPT — see e.g. Hindmarsh:2020hop for a review. Motivated by their possible role in NS mergers, in this paper we initiate the study of relativistic, charged, superheated bubbles, focusing on their hydrodynamic properties in the simple case of a bag-model, conformal EoS. We find two qualitative differences compared to supercooled bubbles. First, the pressure inside a superheated bubble may be higher or lower than the pressure outside the bubble. In contrast, in supercooled bubbles the inside pressure is always higher than the outside pressure. By “inside” and “outside” we do not mean “right behind” and “right in front” of the bubble wall, but deep inside the bubble (in the stable phase) and asymptotically far away from the bubble (in the metastable phase) — see Fig. 6(top). The fact that hydrodynamics allows this pressure difference to take either sign in superheated bubbles is remarkable because it shows that the usual intuition, according to which the bubble expands because the pressure inside wins over the combination of outside pressure plus surface tension, is overly simplistic. The second difference is that some fluid flows develop regions behind the bubble wall that are necessarily metastable which would therefore decay at sufficiently long times. As we will explain, this is not true for supercooled transitions.
@@
Part of the contents of this paper overlap with those in Ref. Barni:2024lkj , which appeared while this paper was being typewritten. To facilitate comparison with that reference, we note that what we call deflagrations, hybrids and detonations are dubbed inverse detonations, inverse hybrids and inverse deflagrations in Barni:2024lkj .
II Self-similar profiles
@@
$w\,\partial_{\mu}u^{\mu}+u^{\mu}\,\partial_{\mu}e=0$ $\,$, (3)
$w\,\bar{u}^{\nu}u^{\mu}\,\partial_{\mu}u_{\nu}-\bar{u}^{\mu}\,\partial_{\mu}p=0$ $\,$,
$n\,\partial_{\mu}u^{\mu}+u^{\mu}\,\partial_{\mu}n=0$ $\,.$
@@
In order to obtain a closed system of equations, we must specify a relation between pressure, energy density and charge density, $p(e,n)$. This is sometimes referred to as the “reduced EoS” (see e.g. Komoltsev:2021jzg ) because, although it determines the hydrodynamic flows, we will see in Sec. III that it contains less information than the true EoS, $p(T,\mu)$. If the pressure in each phase obeys $p=e/d$, then the speed of sound is constant, $c_{s}^{2}=1/d$. In this case, (8) becomes a nested system that can be sequentially solved for $v$, $w$ and $n$, similarly to what happens in the neutral case Espinosa:2010hh . The nested structure means that the addition of charge does not change the properties of self-similar flows for conformal EoS. In fact, the last equation simply implies that $n\propto w^{1/(1+c_{s}^{2})}$. In contrast, if the speed of sound is not constant, $c_{s}=c_{s}(e,n)$, then equations (8) are coupled to one another. In this case the effect of the charge is already present at the level of the self-similar flow. We leave studies of non-conformal theories for future work.
@@
where ++(-) refers to the plasma in front (behind) the discontinuity, the velocities $v_{\pm}$ are measured in the rest frame of the discontinuity, and $\gamma$ is the usual Lorentz factor. Rearranging the first two matching conditions we arrive at
@@
Eqs. (8) and (12) determine the entire fluid flow as a function of three input parameters: the wall velocity $\xi_{w}$, the nucleation energy density $e_{n}$ and the nucleation charge density $n_{n}$.
In the following we will solve these equations in the case of spherical bubbles in three spatial dimensions ($d=3$) and a constant speed of sound.
When $c_{s}$ is constant we can solve for the velocity in (8) independently of the rest of the variables. By assuming different initial conditions $v(\xi_{0})=v_{0}$, one can obtain all possible flows that are self-similar, shown in Fig. 3. Contrary to the supercooled case, we seek solutions where at $\xi>0$ (expanding bubbles) the flow has negative speed $v<0$ (energy inflow). In order to produce Fig. 3, we chose $c_{s}^{2}=1/3$. We expect that different values of the speed of sound will change the picture quantitatively but not qualitatively.
@@
There is yet another kind of flows, those that are contained in the shaded area of Fig. 3. Given that they would become multivalued, these flows cannot cross the dot-dashed line, and they must follow to higher $\xi$. As they cannot smoothly connect to a state with $v=0$, they jump discontinuously to it at the shock location. This shock is different to the one present in hybrids and detonations as the flow in motion is behind and not ahead. An entropic analysis shows that the shock front in Fig. 3 increases entropy only if the flow in motion is ahead. Hence, no flow starting in the mentioned area is physically realizable.
@@
III Solutions for a conformal equation of state
Hereafter we focus on the case of three spatial dimensions, $d=3$. In order to analyze the fluid flows in more detail, a choice of an EoS is needed.
Assuming two branches such that $c_{s}^{2}=1/3$ in each of them, we choose
$p_{H}(T,\mu)$ $=a_{H}T^{4}+b_{H}T^{2}\mu^{2}+c_{H}\mu^{4}-\epsilon\,$, (13)
$p_{L}(T,\mu)$ $=a_{L}T^{4}+b_{L}T^{2}\mu^{2}+c_{L}\mu^{4}\,$,
@@ III Solutions for a conformal equation of state
where $a_{H,L}$, $b_{H,L}$ and $c_{H,L}$ reflect the number of degrees of freedom in each phase and “HH” and “LL” refer to the high- and the low-energy phases, respectively. The constant $\epsilon>0$ is sometimes referred to as the “bag constant” and it measures the energy difference between the two phases in vacuum, that is, at $T=\mu=0$. This is easily verified by means of the usual thermodynamic relations
$e+p=Ts+\mu n\,,\qquad s=\frac{\partial P}{\partial T}\,,\qquad n=\frac{\partial P}{\partial\mu}\,$, (14)
which for the two branches in (13) yield
$e_{H}=3p_{H}+\epsilon\,,\qquad e_{L}=3p_{L}\,.$ (15)
Eqs. (15) are what we referred to above as the “reduced EoS”. Note that different microscopic theories give rise to different EoS through different values of the constants $a_{H,L}$, $b_{H,L}$ and $c_{H,L}$, but all EoS of the form (13) lead to the same reduced EoS (15).
The pressure or the energy density are defined up to an overall, unphysical, additive constant $p_{0}$. In other words, in the absence of gravity the physics is invariant under the simultaneous shift [^2] In the presence of gravity, e.g. in Cosmology, $p_{0}$ becomes physical since it affects the expansion rate of the Universe.
$p_{H}\to p_{H}+p_{0}\,,\qquad p_{L}\to p_{L}+p_{0}\,.$ (16)
Here we have fixed this ambiguity without loss of generality by setting
$p_{0}=0\,.$ (17)
This is equivalent to requiring that the pressure vanish in the vacuum state, namely that
$p_{L}(T=0,\mu=0)=0\,.$ (18)
With this convention, the global stability of the vacuum state constraints the parameters $a_{H,L}$, $b_{H,L}$ and $c_{H,L}$ that specify the EoS so that
$p_{L}\geq 0\,.$ (19)
In contrast to $p_{0}$, the bag constant is physical because it measures the pressure or energy difference between the two phases. This introduces a preferred scale that breaks conformal invariance, as reflected for example by the fact the trace of stress tensor, e3pe-3p, is non-zero in the $H$-phase. This scale sets the overall magnitude of the critical temperature and chemical potential at which the phase transition between the two branches takes place, which is determined by the condition
$p_{L}(T_{c},\mu_{c})=p_{H}(T_{c},\mu_{c})\,.$ (20)
We thus see that referring to (13) as a conformal equation of state, as is common in the literature, is an abuse of language. The reason for this terminology is the fact that the speed of sound (9) is 1/3 in both branches, and that the bag constant cancels out in the enthalpy $w=e+p$. In the hydrodynamic context, since only $c_{s}$ and $w$ enter the equations (8), this means that the flows on both sides of the bubble wall are locally the same as in a conformal theory. Conformality is broken at the bubble wall, where the jump implied by the matching conditions (11)-(12) does depend on the bag constant. The entire flow is thus governed solely by the reduced EoS. As a consequence, microscopic theories with different values of the constants $a_{H,L}$, $b_{H,L}$ and $c_{H,L}$ give rise to the same set of flows. However, the stability of these flows does depend on these constants, since they determine the location of the phase transition and hence of the boundaries between stable, metastable and unstable regions, as well as determining if entropy increases or not during the expansion. Therefore, a given flow may be stable according to one EoS and unstable according to another. We will come back to this point below.
In this paper we will choose
$a_{H}=b_{H}=c_{H}=1\,,\qquad a_{L}=b_{L}=c_{L}=1/2\,$, (21)
for which the phase diagram is shown in Fig. 4. There is a curve of FOPT transition at $T_{c}^{2}+\mu_{c}^{2}=\sqrt{2\epsilon}$. Indeed, recall that the thermodynamically preferred phase is the one with the highest pressure (equivalently, the lowest free energy), and we have
$p_{L}>p_{H}\,\,\,\,\,\mbox{if}\,\,\,\,\,T_{c}^{2}+\mu_{c}^{2}<\sqrt{2\epsilon}\,,\qquad p_{H}>p_{L}\,\,\,\,\,\mbox{if}\,\,\,\,\,T_{c}^{2}+\mu_{c}^{2}>\sqrt{2\epsilon}\,.$ (22)
We do not choose this EoS because we expect it to provide a realistic description of QCD but because it is the simplest example in which an investigation of superheated bubbles is possible. We will report on more realistic examples elsewhere.
@@
Figure 4: Phase diagram for the choice $a_{H}=b_{H}=c_{H}=1$ and $a_{L}=b_{L}=c_{L}=1/2$. “LL” and “HH” refer to the low- and high-energy phases, respectively. All quantities are measured in units of the bag constant, $\epsilon$, which is the energy difference between the two phases at $T=\mu=0$.
@@
The specific detonation shown in Fig. 6(middle) has been chosen to illustrate that, for some flows, the pressure in a region just behind the wall, in this case in the region $\xi\lesssim 0.8$, can be negative. This is necessarily lower than the pressure of the $L$-phase regardless of the values of the parameters $a_{H,L}$, $b_{H,L}$ and $c_{H,L}$ that specify the EoS — see Eqn. (19). [^3] Recall that we fixed the ambiguity in the overall value of the pressure by setting $p_{0}$ — see (17). If we had chosen a different value of $p_{0}$, then we would have found that the pressure right behind the wall in Fig. 6(middle) is lower than $p_{0}$, and we would also have concluded that the preferred phase is the $L$-phase. Since the thermodynamically preferred phase is the one with the highest pressure (equivalently, with the lowest free energy), it follows that right behind the wall the preferred phase is the $L$-phase. This can also be understood in terms of the temperature and the chemical potential. Since $p<0$, the state right behind the wall is in the $H$-phase and $T,\mu$ must obey $T^{2}+\mu^{2}<\sqrt{2\epsilon}$. Looking at (22) we conclude again that, under these conditions, the preferred phase is indeed the $L$-phase.
In summary, despite the fact that the flow in Fig. 6(middle) can be realized in many different microscopic theories, it is unstable in all of them. This is in stark contrast with supercooled flows, for which there is always at least one choice of the parameters $a_{H,L}$, $b_{H,L}$ and $c_{H,L}$ for which the flow is stable (although it may be unstable for other choices). This is illustrated, for example, by the results in Caprini:2011uz . Fig. 4 of this reference shows the comparison between $T$ and $T_{c}$ right behind the wall as a function of $a_{-}/a_{+}$ ($a_{L}/a_{H}$ in our terminology). For certain values of this ratio the flow behind the wall is unstable, for others it is stable.
We conclude that, in the superheated case, at arbitrarily late times there is an arbitrarily large, approximately homogeneous region behind the wall that is necessarily thermodynamically disfavoured with respect to the $L$-phase at the same temperature and chemical potential. This region will eventually decay by nucleating bubbles of the $L$-phase if the bubble is allowed to expand for arbitrarily long times. Whether this happens in practice will depend on the nucleation rates in front and behind the bubble wall, on the bubble wall velocities, etc. Nevertheless, from the viewpoint of the dynamics of a single bubble, flows with negative pressure may regarded as thermodynamically unstable. In Fig. 7 these flows correspond to points above the blue curve labelled as $p_{-}=0$. Note that deflagrations cannot suffer from this problem since the phase right behind the wall is directly the stable $H$-phase.
@@ IV Bounds
Let us start with detonations. Given a wall velocity $\xi_{w}$, consider increasing the strength $\alpha_{n}$ or, equivalently, decreasing the nucleation energy $e_{n}$ relative to the bag constant $\epsilon$. This decreases the energy behind the wall $e_{-}$ and increases the speed $v_{-}$.
Clearly, the strongest transition we can consider is that for which the energy density in the fluid behind the wall reaches its vacuum energy, that is, $e_{-}=\epsilon$. Precisely at this point we also reach the maximum possible velocity behind the wall, $v_{-}=1$. Solving the matching conditions with these values results in the bound
@@
We see that, as we increase the strength factor, the first solutions that cease to exist are the hybrid flows. The next ones to be forbidden are detonations. Therefore, for strong transitions only deflagrations survive, but they are confined to regions with very low wall velocities.
This picture is opposite to that in the supercooled case (Espinosa:2010hh, ), where deflagrations cease to exist for strong transitions while hybrids and detonations become confined to higher and higher wall velocities.
@@
V Efficiency factor
Ref. Casalderrey-Solana:2022rrn showed that the dynamics of superheated bubbles in NS mergers can give rise to an interesting GW signal in the MHz range. A relevant quantity to estimate the amount of GW production is the so called efficiency factor Hindmarsh:2020hop , which is a measure of how much energy has been deposited in the motion of the fluid. The energy fraction can always be computed as the fraction of kinetic energy to the total contained in the volume of the the expanding bubble,
$\kappa=\frac{\int_{V_{\text{bubble}}}d^{3}xvT^{tr}}{\int_{V_{\text{bubble}}}d^{3}xT^{tt}}=\frac{\int_{0}^{\xi_{\text{max}}}d\xi\xi^{2}w\gamma^{2}v^{2}}{\int_{0}^{\xi_{\text{max}}}d\xi\xi^{2}\left(w\gamma^{2}-p\right)}=\frac{3\alpha_{n}}{\epsilon\xi_{\text{max}}^{3}}\int_{0}^{\xi_{\text{max}}}d\xi\xi^{2}w\gamma^{2}v.$ (30)
To obtain the shown expression we have used the fact that the volume of the bubble is $V_{\text{bubble}}=4/3\pi\xi_{\text{max}}^{3}t^{3}$, with $\xi_{\text{max}}=c_{s}$ ($\xi_{\text{max}}=v_{w}$) for deflagrations and hybrids (detonations), and replaced the integral in the denominator by the total energy in the original state, before nucleation.
We have computed the efficiency factor for several strength factors as a function of the wall velocity in the allowed regions where we can find bubble solutions. The result is summarized in Fig. 8. The empty regions above the gray dashed line arise from the disallowed regions, above the solid black curves, in Fig. 7.
Figure 8: Efficiency factor (30) as a function of the bubble wall velocity for several choices of the strength factor. The dashed gray line represents the line beyond which no solution is allowed. The dotted gray line represents the boundary between deflagrations and hybrids, while the solid vertical line lies at the speed of sound, where hybrids transition into detonations.
The behavior of the efficiency factor for deflagrations and detonations is quite simple, since it is monotonous in both cases. For deflagrations the efficiency factor increases with the wall velocity at fixed strength, whereas for detonations it decreases. In both cases the efficiency factor increases with the strength factor at fixed wall velocity (until one hits the maximum possible $\kappa$ in dashed gray).
Hybrids exhibit a richer, non-monotonic behaviour. For low $\alpha_{n}$, hybrids connect the deflagrations (along the dotted gray line) with the detonations (at the vertical $v_{w}=c_{s}$ solid line) and the efficiency factor has a maximum at some intermediate value of the wall velocity. As the envelope – in dashed gray – shows, the largest efficiency factor value is found for detonations, increasing with the wall speed up to a maximum value slightly below $0.5$.